1-5. Consider the following probability distribution for the number of pick pockets at the corner of Stolen Street and Robbed Road in Madrid, Spain on the average afternoon. Pick Pockets Proability 0 1 2 3 4 0.0 10 0.2 0.3 0y 1. What percentage of the time are there more than 2 picket pockets? 2. What is the most likely number of pick pockets to be present? 3. What is the avernge number ofpick pocketson the comer? 4. What is the standard deviation of the number of pick pockets? 5. What is the average number of pick pockets in a 30 day month?
Answer #1

1)

P(X>2) = P(X=3)+P(X=4) = 0.3+0.4 = 0.7

2)

Most likely is 4, as the probability is highest for 4.

3) ,4),5)

X P(X) X*P(X) X^2*P(X)
0 0.000 0.000 0.000
1 0.100 0.100 0.100
2 0.200 0.400 0.800
3 0.300 0.900 2.700
4 0.400 1.600 6.400
1.000 3.000 10.000
Mean ? = ? X P(X) 3.000
Variance ? X^2 P(X)- ?^2 1.000
SD SQRT(Var) 1.000

5)

E(Y) = 30*E(X) = 30*3 = 90